OpenStudy (anonymous):

1. Algebraically transform r = -8sinθ into a Cartesian equation of a circle (in standard form).

OpenStudy (amistre64):

algebraically eh.....

OpenStudy (amistre64):

r = -8sin(t) is a straight line right?

OpenStudy (anonymous):

no, it's a circle

OpenStudy (amistre64):

I see it :)...polars still get me a little lol

OpenStudy (anonymous):

Ha same here

OpenStudy (amistre64):

lets do I table to see where it going ...can we do that? t = 0,45,90,360....

OpenStudy (amistre64):

when its at 90; or straight up, we get r = -8 so this is a circle about the y axis with a radius of 8 centerd 4 below the origin...

OpenStudy (amistre64):

but I am prolly messing that up still....

OpenStudy (amistre64):

nah..im right lol

OpenStudy (anonymous):

well in my textbook theres a question similar to this, but i dont know how to get the last equation. r=6sinθ r²=6rsinθ x²y²=6y x²-6y+9+x²=9 (y-3)²+x²=3²

OpenStudy (amistre64):

yeah, I was gettin to that :) just wanted to make sure I had an answer key to go by ;)

OpenStudy (anonymous):

OOHHH sorry

OpenStudy (amistre64):

our equation is going to end up as: x^2 + (y+4)^2 = 16 r = -8sin(t)

OpenStudy (amistre64):

r = -8sin(t) ; muliply by r r^2 = -8rsin(t) ; convert to cartesians x^2 + y^2 = -8y ; +8y x^2 + y^2 +8 = 0 ; complete square for "y" x^2 + y^2 +8y +16 = 16 ; convert to circle equation x^2 + (y+4)^2 = 16 ; tada!!!

OpenStudy (amistre64):

you see thats a typo ...8y*

OpenStudy (anonymous):

At which step? and where did the 16 come from?

OpenStudy (amistre64):

by now you should be familiar with a process called "completing the square". remember it?

OpenStudy (amistre64):

step 4 I put in 8 instead of 8y ;)

OpenStudy (anonymous):

Oh ok I see the typo

OpenStudy (amistre64):

a complete square can be transformed back and forth between to ...... forms.

OpenStudy (amistre64):

(x+4)^2 = x^2 +8x +16 right? we can easily move back and forth between these equations since the are equal to each other...

OpenStudy (amistre64):

but what is usually missing in order to "complete" a sqaure is this: (x+___)^2 = x^2 +8x + ______

OpenStudy (amistre64):

we need 2 numbers that are exactly the same, that add to get 8. what numbers are they?

OpenStudy (anonymous):

4

OpenStudy (amistre64):

good :) 4+4 = 8, that will satisfy that middle term; now we multiply 4*4 to get the last term. 4^2 = 16 (x+4)^2 = x^2 +8x + 16 you see where we got it now?

OpenStudy (anonymous):

yeah

OpenStudy (amistre64):

x^2 + 6x + ____ what would we complete the square with here?

OpenStudy (anonymous):

(x+3)²=x²+6x+12?

OpenStudy (amistre64):

real close....look again. 3+3 = 6 3*3 != 12

OpenStudy (amistre64):

3*3 = ?

OpenStudy (anonymous):

9

OpenStudy (amistre64):

good :) we need to add 9 to "complete" the square what happens when you add a number to one side of an equation? what do we do to the other side?

OpenStudy (anonymous):

Add it on the other side

OpenStudy (amistre64):

Exactly :) Do you see where I did that in the problem? x^2 + y^2 +8 = 0 ; complete square for "y" x^2 + y^2 +8y +16 = 16

OpenStudy (amistre64):

i really should fix that typo lol

OpenStudy (anonymous):

Haha no. That's where I'm stuck.

OpenStudy (amistre64):

x^2 + [ y^2 +8y +____] = 0 ; complete square for "y" what number do we use to "complete" the square for y?

OpenStudy (anonymous):

16

OpenStudy (amistre64):

its always gonna be half the middle term and then square it

OpenStudy (amistre64):

8/2 = 4 -> 4^2 = 16 :) good

OpenStudy (amistre64):

x^2 + [y^2 +8y +16] = 0 + 16 ^^^ ^^^ we add 16 to both sides righ there right?

OpenStudy (anonymous):

Yeah

OpenStudy (amistre64):

then we just clean it up: x^2 + (y+4)^2 = 16

OpenStudy (amistre64):

which agrees with what we drew in the first place ;)

OpenStudy (anonymous):

Where did the 8y go then?

OpenStudy (amistre64):

remember we can move between the forms of a complete square? (y+4)^2 = y^2 +8y +16 we just use the one for the other.... they are identical in value, they only look different in form.

OpenStudy (anonymous):

Ohhh ok. I get it now!!

OpenStudy (amistre64):

once we got a "complete" square, we use it to clean up the equation ;)

OpenStudy (anonymous):

Ok, I finally understand this now! Thanks for helping me out!

OpenStudy (amistre64):

youre welcome :)